Udzungwa (01.01.2013 to 03.31.2013)

You are here

PDF versionPDF version
Year: 
2013
Quarter: 
January - March
Communications activities: 

none

Unusual Events at the TEAM Site: 

none

New Species: 

none

Protocol Activities: 

During this period, only Climate protocol was being implemented. Two sensors namely HMP45AC and Li200 that were sent for calibration earlier on have been received and fixed back into the datalogger. In addition, another set of sensors namely HMP45AC and Li200 have been removed ready for calibration.

Protocol Problems: 

none

Schedule Problems: 

none

Logistical Problems: 

none

Suggestions: 

none